# Optimizing the pump coupling for a three-wave mixing Josephson parametric amplifier

### v1

@@ VI Conclusion and discussion
In this work we have demonstrated a 3 orders of magnitude improvement in both power efficiency and pump leakage suppression on a filter-coupled SNAIL parametric amplifier compared to a regular capacitor-coupled counterpart. This improvement is achieved by implementing 3-stage, on-chip filters on the signal and pump ports of the device. Most previous work incorporating microwave filters into JPA design were targeted at improving the bandwidth and dynamic range[5, 9, 11, 12]. Our work addresses the increasingly important issue of pump power delivery for 3-wave-mixing JPAs. While a recent study [12] discussed the use of an external diplexer to facilitate pump delivery, our work provides the methodology for optimizing the pump coupling at the device design level. The on-chip filtering strategy offers advantages in terms of compactness and eliminates insertion loss from additional external components. Moreover, the network-based description we introduced, particularly the formulation of the resonance condition in terms of Thevenin impedance, provides a systematic framework for designing the embedding circuit for Josephson elements with frequency-dependent coupling to the environment. We believe the strategy of including microwave filters in circuit QED design should benefit many other applications featuring off-resonant parametric drives[33, 34, 35, 36, 37].
Moreover, our work shows the benefits of the signal-pump frequency separation of 3-wave mixing amplifiers, which allows for substituting pump line attenuation with filtering, reducing both the cryogenic power dissipation and the overall pump power required from the room-temperature rf generator. In particular, the filter-coupled SPA is robust against thermal noise on the pump port with up to 4 K noise temperature. This demonstration opens up the possibility for eliminating attenuation at the millikelvin stage of the dilution refrigerator, and for introducing a separate port for pumping kinetic-inductance-based parametric amplifiers that can operate at higher temperature[38]. Such optimizations will greatly alleviate the challenge of operating a large array of quantum-limited amplifiers in a large-scale superconducting quantum processor.
@@ VII Acknowledgement
The authors want to thank José Aumentado for providing the shot noise tunnel junction used for noise calibration. We acknowledge Nicholas Frattini and Benjamin Chapman for helpful discussions. W.D. thanks Luigi Frunzio, Robert Schoelkopf, and A. Douglas Stone for their advice on this project. This material is based upon work supported by the U.S. Department of Energy, Office of Science, National Quantum Information Science Research Centers, Co-design Center for Quantum Advantage (C2QA) under contract number DE-SC0012704. S.S. acknowledges support from the Air Force Office of Scientific Research (grant numbers FA9550-20-1-0177, FA9550-22-1-0203) and the Army Research Office (grant numbers W911NF-23-1-0096, W911NF-23-1-0251). Fabrication facilities use was supported by the Yale Institute for Nanoscience and Quantum Engineering (YINQE) and the Yale SEAS Cleanroom. The views and conclusions contained in this document are those of the authors and should not be interpreted as representing the official policies, either expressed or implied, of the US Government. The US Government is authorized to reproduce and distribute reprints for Government purposes notwithstanding any copyright notation herein.
@@ Appendix A Device fabrication and measurement setup
The fabrication and packaging processes of our device are originally developed and introduced in Ref. [8, 15]. The aluminum strip (bright pattern in Fig. 3(b)) along with an array of $M=10$ SNAILs embedded inline is fabricated with a single-step e-beam lithography and liftoff process. Each SNAIL consists of three large Josephson junctions with critical current $I_{c}=10.1$ µA in parallel with one small junction with critical current $I_{c}=0.8$ µA, all formed by the Dolan bridge process, as shown in the SEM image in Fig. 3(b) inset.
@@
The FSPA and CSPA were measured with the same wiring setup in two consecutive cooldowns, and in a third cooldown we used a Shot-Noise Tunnel Junction (SNTJ) [39] to calibrate the output chain. When measuring the CSPA device, a lowpass filter was added to the output chain in order to prevent the pump leakage from saturating the room temperature MITEQ amplifier. The two output line setups, with and without the lowpass filter, are calibrated respectively with the SNTJ.
@@ A.1 Noise calibration
The SNTJ serves as a self-calibrated noise source [31] with which we can calibrate the gain $G_{\mathrm{sys}}$ and noise temperature $T_{\mathrm{sys}}$ of the output chain, shown in Figure.8. The uncertainty for noise temperature calibration is dominated by the impedance mismatch on the SNTJ and loss in the bias-tee, and we accordingly extract the errorbar as a function of frequency.
@@ B.1 Scattering analysis for the 3-port network
$=\frac{S_{31}}{1-\Gamma_{3}S_{33}}V^{\mathrm{in},1}+\frac{S_{32}}{1-\Gamma_{3}S_{33}}V^{\mathrm{in},2}$
from which we can obtain the relation in equation (4) considering the pump applied only from port 2 (i.e. $V^{\mathrm{in},1}_{p}=0$).
The signal, on the other hand, is applied only from port 1 (i.e. $V^{\mathrm{in},2}_{s}=0$). Similar expression for $V^{\mathrm{out},1}$ and $V^{\mathrm{out},2}$ can be evaluated accordingly, that gives rise to equation (6) and (8) respectively.
@@ B.2 Signal flow graph analysis for concatenating two-port networks
The standard filter designs typically aim to optimize the insertion loss $|t|$ and return loss $|r|$, assuming both ports are $Z_{0}$ matched. However, in our device the networks combine in the following manner: the Josephson dipole connects a pair of terminals, and the other pair of terminals are galvanically connected via the microstrip ground. Signal-flow graph analysis[29] can be applied to Fig. 10(c) to construct the effective 3-port scattering matrix for the concatenated circuit.
@@ Appendix C Resonance condition from Thevenin circuit method
In this appendix, we formulate the resonance condition for a general circuit containing a lumped Josephson dipole, such as the example shown in Fig. 10(a). Our method can efficiently extract the resonance frequency, damping rate, and Josephson participation ratio [40] of a classical mode, whose Kerr nonlinearity is much smaller than its damping rate. We expect our method to provide some insights into designing Josephson-element-based parametric amplifiers and couplers with complicated embedding circuits.
@@
For obtaining the linear eigenmode of the system, we replace the nonlinear Josephson dipole by its inductance $L_{J}$, with a linear response: $V^{\mathrm{J}}=L_{J}\frac{\mathrm{d}{I^{\mathrm{J}}}}{\mathrm{d}{t}}$. To find the ‘undriven’ modes of motion allowed by a homogeneous linear system, a Laplace transform (represented by $\mathcal{L}$) can be applied to the equation of motion [41]:
$sL_{J}\mathcal{L}\{I^{\mathrm{J}}(t)\}+Z^{\mathrm{Th}}(s)\mathcal{L}\{I^{\mathrm{J}}(t)\}=0$ (18)
@@
Plugging in the resonance condition Eq. 5 in the main text, i.e. $Z^{\mathrm{Th}}(j\omega_{a}-\frac{\kappa_{a}}{2})=-j\omega_{a}L_{J}+\frac{\kappa_{a}}{2}L_{J}$:
$\mathrm{Im}Z^{\mathrm{Th}}[\omega_{a}]$ $=-\left(\omega_{a}L_{J}+\frac{\kappa_{a}}{2}\mathrm{Re}Z^{\mathrm{Th}^{\prime}}[\omega_{a}]\right)$ (22a)
$\mathrm{Re}Z^{\mathrm{Th}}[\omega_{a}]$ $=\frac{\kappa_{a}}{2}\left(L_{J}+\mathrm{Im}Z^{\mathrm{Th}^{\prime}}[\omega_{a}]\right)$ (22b)
@@
Figure 11: Simulated output impedance towards signal port $Z^{(s)}$ and output impedance towards pump port $Z^{(p)}$. The resonance frequency is determined by the point where the black curve (an example with $L_{J}(\Phi_{\mathrm{ext}})$ evaluated at $\Phi_{\mathrm{ext}}=0.3\Phi_{0}$) crosses 0. The slope at this point is related to the Josephson participation of the mode. Real part of $Z^{(s)}$ evaluated at the resonance frequency determines the linewidth of the mode.
In Fig. 11, we show the $Z^{(s)}$ and $Z^{(p)}$ simulated (using AWR Axiem) around the desired resonance frequency. We also plot the imaginary part of $Z^{(s)}[\omega]+Z^{(p)}[\omega]+j\omega L_{J}(\Phi_{\mathrm{ext}})$ evaluated at $\Phi_{\mathrm{ext}}=0.3\Phi_{0}$ as the black curve. As the second term on the right hand side of Eq.(22a) is negligible in our case, the resonance frequency $\omega_{a}$ is essentially the point where the black curve crosses 0. We then evaluate $\kappa_{a}$ from Eq.(22b), using $\mathrm{Re}Z^{\mathrm{Th}}$ and $\mathrm{Im}Z^{\mathrm{Th}^{\prime}}$ taken at $\omega_{a}$. Since the real part of $Z^{(p)}$ is almost 0 around resonance frequency (as the pump-port filter rejects the on-resonance signal), the damping rate of the mode originates entirely from the dissipation at signal port.
In Fig. 3(c) from the main text we have shown the resonance frequency and linewidth extracted using the method above, with varying $L_{J}(\Phi_{\mathrm{ext}})$ for the SNAIL array[28]. They agree with the experimental results fairly well.
Lastly, we can also extract the Josephson inductive participation ratio of each mode $a$ from the Thevenin circuit method. Following the same ideas as blackbox circuit quantization[42], the Josephson dipole can be expressed as a linear inductor $L_{J}$ in parallel with a purely nonlinear element. Under the assumption of weak dissipation (i.e. $\kappa_{a}\ll\omega_{a}$) for all eigenmodes of interest, the full linear circuit consisting of $L_{J}$ in parallel with the Thevenin impedance can be decomposed into a set of poles in the form of parallel RLCs, i.e. an approximate Foster-equivalent admittance.
$\frac{1}{sL_{J}}+\frac{1}{Z^{\mathrm{Th}}(s)}=\sum_{a}\left(sC_{a}+\frac{1}{sL_{a}}+\frac{1}{R_{a}}\right)$ (25)
@@
$p_{a}^{\mathrm{J}}:=\frac{L_{a}}{L_{J}}\approx\frac{2}{1+\frac{\mathrm{Im}Z^{\mathrm{Th}^{\prime}}[\omega_{a}]}{L_{J}}}$ (26)
which by definition is the inductive participation ratio (twice the energy participation ratio as defined in [40]) of the Josephson dipole in mode $a$.
An observation is made that $p_{a}^{\mathrm{J}}$ enters equation (22b) such that $\mathrm{Re}Z^{\mathrm{Th}}[\omega_{a}]/L_{J}=\kappa_{a}/p_{a}^{\mathrm{J}}$. It can be understood as a scaling factor between the damping rate at the Josephson dipole and that for the the resonance mode.
@@ Appendix D Modeling the pumped Josephson dipole
In this appendix, we introduce a circuit model describing a current-pumped parametric process in a single lumped Josephson dipole, in analogy with previous work on flux-pumped parametric processes [26]. This description applies to the synthesis of parametric amplifiers and converters, as well as of the current-pumped ancilla for gaussian operations on bosonic modes [43, 36, 44].
Similar to the Taylor expansion of the Josephson relation $I=\frac{\phi_{0}}{L_{J}}\sin\varphi$ for a single Josephson junction, the electric current $I^{\mathrm{J}}(t)$ is linked to the superconducting phase $\varphi^{\mathrm{J}}(t)=\Phi^{\mathrm{J}}(t)/\phi_{0}$ across an arbitrary Josephson dipoles (such as the SNAIL[28], rf SQUID or Gradiometric-SNAIL[45]) by the following relation:
$I^{\mathrm{J}}=\frac{\phi_{0}}{L_{J}}\left(\varphi^{\mathrm{J}}+\frac{c_{3}}{2}\varphi^{\mathrm{J}2}+\cdots\right)$ (27)
Here, we keep only the 3rd order of nonlinearity and study the parametric squeezing processes arising from this term.
@@
$I^{\mathrm{J}}_{s}$ $=\frac{\phi_{0}}{L_{J}}\left(c_{2}\varphi^{\mathrm{J}}_{s}+\frac{c_{3}}{2}\varphi^{\mathrm{J}}_{p}\varphi^{\mathrm{J}*}_{i}\right)$ (29)
$I^{\mathrm{J}*}_{i}$ $=\frac{\phi_{0}}{L_{J}}\left(c_{2}\varphi^{\mathrm{J}*}_{i}+\frac{c_{3}}{2}\varphi^{\mathrm{J}*}_{p}\varphi^{\mathrm{J}}_{s}\right)$
The stiff pump approximation assumes that the existence of probes at other frequencies does not affect the steady state $V^{\mathrm{J}}_{\omega_{p}}$ evaluated from linear circuit analysis. Under such an approximation, the harmonic component $\varphi^{\mathrm{J}}_{p}=\frac{V^{\mathrm{J}}_{p}}{j\omega_{p}\phi_{0}}$ can be treated as a parameter (instead of a variable). The harmonic balance equation (29) between $\omega_{s}$ and $\omega_{i}$ harmonic components is thus linearizable, which can be written as a cross-admittance matrix:
$\left(\begin{matrix}I^{\mathrm{J}}_{s}\\ I^{\mathrm{J}*}_{i}\end{matrix}\right)=\frac{1}{jL_{J}}\left(\begin{matrix}\frac{1}{\omega_{s}}&\frac{\frac{c_{3}}{2}\varphi^{\mathrm{J}}_{p}}{-\omega_{i}}\\ \frac{\frac{c_{3}}{2}\varphi^{\mathrm{J}*}_{p}}{\omega_{s}}&\frac{1}{-\omega_{i}}\end{matrix}\right)\left(\begin{matrix}V^{\mathrm{J}}_{s}\\ V^{\mathrm{J}*}_{i}\end{matrix}\right)$ (30)
@@
$Z^{\mathrm{Th}}[\omega_{i}]\approx-j\omega_{i}L_{J}\left(1+j\frac{1}{p^{\mathrm{J}}_{i}Q_{i}}\right)$ (33)
@@
$Y^{\mathrm{nl}}\left(V^{\mathrm{J}}_{p}\right)$ $=\frac{1}{j\omega_{s}L_{J}}\frac{\left|\frac{c_{3}}{2}\varphi^{\mathrm{J}}_{p}\right|^{2}}{\frac{1}{1-j\frac{1}{p^{\mathrm{J}}_{i}Q_{i}}}-1}$ (34)
$\approx-\frac{p^{\mathrm{J}}_{i}Q_{i}}{\omega_{s}L_{J}}\left|\frac{c_{3}}{2}\varphi^{\mathrm{J}}_{p}\right|^{2}$

### v2

@@ VI Conclusion and discussion
In this work, we have demonstrated improvements of 3 orders of magnitude in both power efficiency and pump-leakage suppression on a filter-coupled SNAIL parametric amplifier compared to its conventional capacitor-coupled counterpart. This improvement is achieved by implementing three-stage, on-chip filters on the signal and pump ports of the device. Most previous works incorporating microwave filters into JPA designs have targeted at improving the bandwidth and dynamic range Roy et al. (2015); Naaman et al. (2019); Kaufman et al. (2023, 2024). Our work addresses the increasingly important issue of pump power delivery for three-wave-mixing JPAs. While a recent study  Kaufman et al. (2024) discussed the use of an external diplexer to facilitate pump delivery, our work provides the methodology for optimizing the pump coupling at the device design level. The on-chip filtering strategy offers advantages in terms of compactness and eliminates insertion loss from additional external components. Moreover, the network-based description we introduced, particularly the formulation of the resonance condition in terms of Thevenin impedance, provides a systematic framework for designing the embedding circuit for Josephson elements with frequency-dependent coupling to the environment. We believe the strategy of including microwave filters in circuit QED design should benefit many other applications featuring off-resonant parametric drives Grimm et al. (2020); Campagne-Ibarcq et al. (2020); Zhou et al. (2023); Chapman et al. (2023); Xia et al. (2023).
Moreover, our work shows the benefits of the signal-pump frequency separation of three-wave mixing amplifiers, which allows for substituting pump line attenuation with filtering, reducing both the cryogenic power dissipation and the overall pump power required from the room-temperature rf generator. In particular, the filter-coupled SPA is robust against thermal noise on the pump port with up to 4 K noise temperature. This demonstration opens up the possibility of eliminating attenuation at the millikelvin stage of the dilution refrigerator, and introducing a separate port for pumping kinetic-inductance-based parametric amplifiers that can operate at higher temperatures Xu et al. (2024). Such optimizations will greatly alleviate the challenge of operating a large array of quantum-limited amplifiers in a large-scale superconducting quantum processor.
@@ VII Acknowledgement
The authors would like to thank José Aumentado for providing the shot-noise tunnel junction used for noise calibration. We acknowledge Nicholas Frattini and Benjamin Chapman for helpful discussions. W.D. thanks Luigi Frunzio, Robert Schoelkopf, and A. Douglas Stone for their advice on this project. This material is based upon work supported by the U.S. Department of Energy, Office of Science, National Quantum Information Science Research Centers, Co-design Center for Quantum Advantage (C2QA) under contract number DE-SC0012704. S.S. acknowledges support from the Air Force Office of Scientific Research (Grants No. FA9550-20-1-0177 and No. FA9550-22-1-0203) and the Army Research Office (Grants No. W911NF-23-1-0051, No. W911NF-23-1-0096, and No. W911NF-23-1-0251). The use of fabrication facilities was supported by the Yale Institute for Nanoscience and Quantum Engineering (YINQE) and the Yale SEAS Cleanroom. The views and conclusions contained in this document are those of the authors and should not be interpreted as representing the official policies, either expressed or implied, of the U.S. Government. The U.S. Government is authorized to reproduce and distribute reprints for Government purposes notwithstanding any copyright notation herein.
@@ Appendix A Device fabrication and measurement setup
The fabrication and packaging processes of our device are originally developed and introduced in Ref. Frattini et al. (2018); Sivak et al. (2019). The aluminum strip (bright pattern in Fig. 3(b)) along with an array of $M=10$ SNAILs embedded inline is fabricated with a single-step e-beam lithography and liftoff process. Each SNAIL consists of three large Josephson junctions with critical current $I_{c}=10.1$ µA in parallel with one small junction with critical current $I_{c}=0.8$ µA, all formed by the Dolan bridge process, as shown in the SEM image in Fig. 3(b) inset.
@@
The FSPA and CSPA were measured with the same wiring setup in two consecutive cooldowns, and in a third cooldown we used a Shot-Noise Tunnel Junction (SNTJ) Spietz et al. (2003) to calibrate the output chain. When measuring the CSPA device, a lowpass filter was added to the output chain in order to prevent the pump leakage from saturating the room temperature MITEQ amplifier. The two output line setups, with and without the lowpass filter, are calibrated respectively with the SNTJ.
@@ A.1 Noise calibration
The SNTJ serves as a self-calibrated noise source Malnou et al. (2024) with which we can calibrate the gain $G_{\mathrm{sys}}$ and noise temperature $T_{\mathrm{sys}}$ of the output chain, shown in Figure.8. The uncertainty for noise temperature calibration is dominated by the impedance mismatch on the SNTJ and loss in the bias-tee. We accordingly extract the errorbar as a function of frequency using the same model as in Sivak et al. (2020).
@@ B.1 Scattering analysis for the 3-port network
$=\frac{S_{31}}{1-\Gamma_{3}S_{33}}V^{\mathrm{in},1}+\frac{S_{32}}{1-\Gamma_{3}S_{33}}V^{\mathrm{in},2}.$
Since the pump is applied only from port 2 (i.e. $V^{\mathrm{in},1}_{p}=0$), the pump coupling efficiency from Eq.(3) can be expressed in terms of scattering parameters as:
$\eta_{p}^{\mathrm{c}}=\left|\frac{1+\Gamma_{3}[\omega_{p}]}{1-\Gamma_{3}[\omega_{p}]S_{33}[\omega_{p}]}S_{32}[\omega_{p}]\right|^{2}\frac{Z_{0}}{\omega_{p}L_{\mathrm{J}}}.$ (10)
The signal, on the other hand, is applied only from port 1 (i.e. $V^{\mathrm{in},2}_{s}=0$). Signal power gain is expected in reflection from port 1, which is given by:
where $\Gamma_{3}[\omega_{s}]$ is the reflection coefficient from the Josephson dipole at the signal frequency. In our model, as illustrated in Fig. 2(c), signal amplification translates into a negative real part for the pumped admittance $Y^{\mathrm{nl}}(V^{\mathrm{J}}_{p})$. Furthermore, achieving high gain ($G\gg 1$) requires $1-\Gamma_{3}[\omega_{s}]S_{33}[\omega_{s}]\approx 0$, which is satisfied for near-resonance signals.
Evaluating $V^{\mathrm{out},2}_{s}$ in a similar manner, we can also write the transmission coefficient from port 1 to port 2:
$T=\left|\frac{V^{\mathrm{out},2}_{s}}{V^{\mathrm{in},1}_{s}}\right|^{2}=\left|S_{21}[\omega_{s}]+\frac{S_{23}[\omega_{s}]\Gamma_{3}[\omega_{s}]S_{31}[\omega_{s}]}{1-\Gamma_{3}[\omega_{s}]S_{33}[\omega_{s}]}\right|^{2}.$ (12)
In order to achieve a noise performance approaching the quantum limit, the signal loss via port 2 should be minimized so that $T\ll G$, essentially demanding $\left|S_{23}[\omega_{s}]\right|\ll\left|S_{13}[\omega_{s}]\right|$.
@@ B.2 Signal flow graph analysis for concatenating two-port networks
The standard filter designs typically aim to optimize the insertion loss $|t|$ and return loss $|r|$, assuming both ports are $Z_{0}$ matched. However, in our device the networks combine in the following manner: the Josephson dipole connects a pair of terminals, and the other pair of terminals are galvanically connected via the microstrip ground. Signal-flow graph analysis Pozar (2005) can be applied to Fig. 10(c) to construct the effective 3-port scattering matrix for the concatenated circuit.
@@ Appendix C Resonance condition from Thevenin circuit method
In this appendix, we formulate the resonance condition for a general circuit containing a lumped Josephson dipole, such as the example shown in Fig. 10(a). Our method can efficiently extract the resonance frequency, damping rate, and Josephson participation ratio Minev et al. (2021) of a classical mode, whose Kerr nonlinearity is much smaller than its damping rate. We expect our method to provide some insights into designing Josephson-element-based parametric amplifiers and couplers with complicated embedding circuits.
@@
For obtaining the linear eigenmode of the system, we replace the nonlinear Josephson dipole by its inductance $L_{\mathrm{J}}$, with a linear response: $V^{\mathrm{J}}=L_{\mathrm{J}}\frac{\mathrm{d}{I^{\mathrm{J}}}}{\mathrm{d}{t}}$. To find the ‘undriven’ modes of motion allowed by a homogeneous linear system, a Laplace transform (represented by $\mathcal{L}$) can be applied to the equation of motion Oppenheim et al. (2010):
$sL_{\mathrm{J}}\mathcal{L}\{I^{\mathrm{J}}(t)\}+Z^{\mathrm{Th}}(s)\mathcal{L}\{I^{\mathrm{J}}(t)\}=0$ (19)
@@
Plugging in the resonance condition Eq. 5 in the main text, i.e. $Z^{\mathrm{Th}}(j\omega_{a}-\frac{\kappa_{a}}{2})=-j\omega_{a}L_{\mathrm{J}}+\frac{\kappa_{a}}{2}L_{\mathrm{J}}$:
$\mathrm{Im}Z^{\mathrm{Th}}[\omega_{a}]$ $=-\left(\omega_{a}L_{\mathrm{J}}+\frac{\kappa_{a}}{2}\mathrm{Re}Z^{\mathrm{Th}^{\prime}}[\omega_{a}]\right)$ (23a)
$\mathrm{Re}Z^{\mathrm{Th}}[\omega_{a}]$ $=\frac{\kappa_{a}}{2}\left(L_{\mathrm{J}}+\mathrm{Im}Z^{\mathrm{Th}^{\prime}}[\omega_{a}]\right)$ (23b)
@@
Figure 11: Simulated output impedance towards signal port $Z^{(s)}$ and output impedance towards pump port $Z^{(p)}$. The resonance frequency is determined by the point where the black curve (an example with $L_{\mathrm{J}}(\Phi_{\mathrm{ext}})$ evaluated at $\Phi_{\mathrm{ext}}=0.3\Phi_{0}$) crosses 0. The slope at this point is related to the Josephson participation of the mode. Real part of $Z^{(s)}$ evaluated at the resonance frequency determines the linewidth of the mode.
In Fig. 11, we show the $Z^{(s)}$ and $Z^{(p)}$ simulated (using AWR Axiem) around the desired resonance frequency. We also plot the imaginary part of $Z^{(s)}[\omega]+Z^{(p)}[\omega]+j\omega L_{\mathrm{J}}(\Phi_{\mathrm{ext}})$ evaluated at $\Phi_{\mathrm{ext}}=0.3\Phi_{0}$ as the black curve. As the second term on the right hand side of Eq.(23a) is negligible in our case, the resonance frequency $\omega_{a}$ is essentially the point where the black curve crosses 0. We then evaluate $\kappa_{a}$ from Eq.(23b), using $\mathrm{Re}Z^{\mathrm{Th}}$ and $\mathrm{Im}Z^{\mathrm{Th}^{\prime}}$ taken at $\omega_{a}$. Since the real part of $Z^{(p)}$ is almost 0 around resonance frequency (as the pump-port filter rejects the on-resonance signal), the damping rate of the mode originates entirely from the dissipation at the signal port.
In Fig. 3(c) from the main text we have shown the resonance frequency and linewidth extracted using the method above, with varying $L_{\mathrm{J}}(\Phi_{\mathrm{ext}})$ for the SNAIL arrayFrattini et al. (2017). They agree with the experimental results fairly well.
Lastly, we can also extract the Josephson inductive participation ratio of each mode $a$ from the Thevenin circuit method. Following the same ideas as blackbox circuit quantization Nigg et al. (2012), the Josephson dipole can be expressed as a linear inductor $L_{\mathrm{J}}$ in parallel with a purely nonlinear element. Under the assumption of weak dissipation (i.e. $\kappa_{a}\ll\omega_{a}$) for all eigenmodes of interest, the full linear circuit consisting of $L_{\mathrm{J}}$ in parallel with the Thevenin impedance can be decomposed into a set of poles in the form of parallel RLCs, i.e. an approximate Foster-equivalent admittance.
$\frac{1}{sL_{\mathrm{J}}}+\frac{1}{Z^{\mathrm{Th}}(s)}=\sum_{a}\left(sC_{a}+\frac{1}{sL_{a}}+\frac{1}{R_{a}}\right)$ (26)
@@
$p_{a}^{\mathrm{J}}:=\frac{L_{a}}{L_{\mathrm{J}}}\approx\frac{2}{1+\frac{\mathrm{Im}Z^{\mathrm{Th}^{\prime}}[\omega_{a}]}{L_{\mathrm{J}}}}$ (27)
which by definition is the inductive participation ratio (twice the energy participation ratio as defined in Minev et al. (2021)) of the Josephson dipole in mode $a$.
An observation is made that $p_{a}^{\mathrm{J}}$ enters equation (23b) such that $\mathrm{Re}Z^{\mathrm{Th}}[\omega_{a}]/L_{\mathrm{J}}=\kappa_{a}/p_{a}^{\mathrm{J}}$. It can be understood as a scaling factor between the damping rate at the Josephson dipole and that for the the resonance mode.
@@ Appendix D Modeling the pumped Josephson dipole
In this appendix, we introduce a circuit model describing a current-pumped parametric process in a single lumped Josephson dipole, in analogy with previous work on flux-pumped parametric processes Sundqvist and Delsing (2014). This description applies to the synthesis of parametric amplifiers and converters, as well as of the current-pumped ancilla for gaussian operations on bosonic modes Gao et al. (2019); Chapman et al. (2023); Ding et al. (2024).
Similar to the Taylor expansion of the Josephson relation $I=\frac{\phi_{0}}{L_{\mathrm{J}}}\sin\varphi$ for a single Josephson junction, the electric current $I^{\mathrm{J}}(t)$ is linked to the superconducting phase $\varphi^{\mathrm{J}}(t)=\Phi^{\mathrm{J}}(t)/\phi_{0}$ across an arbitrary Josephson dipoles by the following relation:
$I^{\mathrm{J}}=\frac{\phi_{0}}{L_{\mathrm{J}}}\left(\varphi^{\mathrm{J}}+\frac{c_{3}}{2}\varphi^{\mathrm{J}2}+\cdots\right)$ (28)
where $L_{\mathrm{J}}$ is the total inductance of the Josephson dipole. Derivation of $L_{\mathrm{J}}$ and the expansion coefficient $c_{3}$ depends on the physical realization of the dipole, examples are present for a lumped array of SNAIL Frattini et al. (2017), rf SQUID or Gradiometric-SNAIL Miano et al. (2022). Here, we have kept only the 3rd order nonlinearity and study the parametric squeezing processes arising from this term.
@@
$I^{\mathrm{J}}_{s}$ $=\frac{\phi_{0}}{L_{\mathrm{J}}}\left(\varphi^{\mathrm{J}}_{s}+\frac{c_{3}}{2}\varphi^{\mathrm{J}}_{p}\varphi^{\mathrm{J}*}_{i}\right)$ (30)
$I^{\mathrm{J}*}_{i}$ $=\frac{\phi_{0}}{L_{\mathrm{J}}}\left(\varphi^{\mathrm{J}*}_{i}+\frac{c_{3}}{2}\varphi^{\mathrm{J}*}_{p}\varphi^{\mathrm{J}}_{s}\right)$
The stiff-pump approximation assumes that the existence of probes at other frequencies does not affect the steady state $V^{\mathrm{J}}_{\omega_{p}}$ evaluated from linear circuit analysis. Under such an approximation, the harmonic component $\varphi^{\mathrm{J}}_{p}=\frac{V^{\mathrm{J}}_{p}}{j\omega_{p}\phi_{0}}$ can be treated as a parameter (instead of a variable). The harmonic balance equation (30) between $\omega_{s}$ and $\omega_{i}$ harmonic components is thus linearizable, which can be written as a cross-admittance matrix:
$\left(\begin{matrix}I^{\mathrm{J}}_{s}\\ I^{\mathrm{J}*}_{i}\end{matrix}\right)=\frac{1}{jL_{\mathrm{J}}}\left(\begin{matrix}\frac{1}{\omega_{s}}&\frac{\frac{c_{3}}{2}\varphi^{\mathrm{J}}_{p}}{-\omega_{i}}\\ \frac{\frac{c_{3}}{2}\varphi^{\mathrm{J}*}_{p}}{\omega_{s}}&\frac{1}{-\omega_{i}}\end{matrix}\right)\left(\begin{matrix}V^{\mathrm{J}}_{s}\\ V^{\mathrm{J}*}_{i}\end{matrix}\right)$ (31)
@@
$Z^{\mathrm{Th}}[\omega_{i}]\approx-j\omega_{i}L_{\mathrm{J}}\left(1+j\frac{1}{p^{\mathrm{J}}_{i}Q_{i}}\right)$ (34)
@@
$Y^{\mathrm{nl}}\left(V^{\mathrm{J}}_{p}\right)$ $=\frac{1}{j\omega_{s}L_{\mathrm{J}}}\frac{\left|\frac{c_{3}}{2}\varphi^{\mathrm{J}}_{p}\right|^{2}}{\frac{1}{1-j\frac{1}{p^{\mathrm{J}}_{i}Q_{i}}}-1}$ (35)
$\approx-\frac{p^{\mathrm{J}}_{i}Q_{i}}{\omega_{s}L_{\mathrm{J}}}\left|\frac{c_{3}}{2}\varphi^{\mathrm{J}}_{p}\right|^{2}$